# Do Visual-Language Maps Capture Latent Semantics?

### v1

@@ V Experiments
The main motivation of the experiments is to demonstrate the proposed benchmark. The two main questions the benchmark aims to answer with the experiments are:
@@
    How consistent are their visual-language embeddings within and across maps?
@@ V-A Data set
The data set used in the benchmark is the Matterport3D data set [12], which consists of houses captured with an RGBD camera and ground truth semantic segmentation. The data set was used in the original work of VLMaps, in which they selected a sample of 10 sequences from the data set. In the benchmark, we use the same sequences and the same set of 42 labels forming the set $\mathcal{L}$; their names are also used in the following experiments as the query set $\mathcal{Q}$ to leverage the semantic ground truth provided with the dataset [^2] Supplementary material including the list of labels will be available in arXiv and link provided here upon acceptance of this work..
@@ V-B Parametrization of the methods
Instead of using the 2D mapping method proposed in the original work [8], the VLMaps map was created using the new 3D mapping method available at the repository of the original authors [33]. This choice was made to create a fair comparison using state-of-the-art methods. The central idea is the same, except the 3D map avoids the problems when aggregating the embeddings along the $z$-axis. All of the parameters were the default used by the original authors. The 3D grid voxel size used is $0.05$ m, as proposed by the authors.
@@
When queried with a query $q$, VLMaps creates a list of queries by setting $q$ and string ”other” into a list of phrases such as ”A photo of …” and ”there is … in the scene”. Then, the most similar query is selected from the list for each voxel. The binary mask is $true$ for voxels where the most similar query is a phrase containing $q$; otherwise, it is $false$. The binary mask is then dilated to encompass whole object instances.
OpenScene maps were created using the parameters proposed in the original paper. We use the proposed 2D-3D ensemble features. The 3D grid voxel size used is $0.02$ m, as proposed by the authors. OpenScene does not provide a way to query open-vocabulary binary masks. Because we use the set of labels as the queries, the binary query mask is created by comparing the equality of the voxel labels and the query.
@@
Figure 2: The intra-map consistency ratios of the methods as histograms, with a kernel density estimation, which is smoothed with a Gaussian kernel. The mean of the distribution is depicted with a vertical line of corresponding color.
@@
V-C Results
V-C1 Queryability
The results of the queryability benchmark are presented in Table I, where the F1-score, precision, recall, and IoU of the method and encoder combinations are presented for the voxel- and instance-based tests. Overall, OpenScene performs better than VLMaps with both encoders, based on the higher F1-scores and IoUs. The higher recall shown by VLMaps with LSeg is likely the result of the post-processing of the queries, where the query results are dilated to encompass whole objects. The instance-based results mostly align with the voxel-based results, except for the slightly decreased performance in all metrics.
While OpenScene with LSeg is better in the voxel-based experiment, both encoders have virtually the same performance in the instance-based experiment when used within OpenScene. This aligns with the findings presented in [11], where OpenScene with LSeg encoder was found to have better IoU.
Therefore, these results show that both the mapping method and choice of encoder matter in the map creation process. They also indicate that the 3D structure of the environment, included by OpenScene, contains information that can be leveraged in addition to the purely image-based creation of embeddings.
All methods have better recall than precision, which are usually tradeoffs. Better recall might be beneficial for queryability, especially in a navigation context: having more object candidates to explore than missing potential objects is better. This is especially visible in VLMaps, where the recall is considerably higher than precision. OpenScene has a more balanced approach, reflected in the high F1-Score, which is a good metric for the overall performance of the methods. IoU also conforms to the results, increasing their reliability.
As the instance classification is based on the prediction of the majority label, many objects seem to match the query only with a region representing less than half of the object. In this regard, the query methodology and, subsequently, the experiment setup could be improved: instead of classifying the results with binary masks, the minimum cosine distance could be used. This could allow the detection of instances with a small but significant match to the query. This would be relevant, for instance, when the object is partially occluded or unobserved. For example, if a sofa is partially covered with a blanket, the image-based methods could segment it as a blanket, but with the visible small region, it could still be recognized as a sofa. In this regard, the 3D understanding could prove extremely valuable: The geometry of the sofa might not significantly change when it is covered, so it can still be recognized as a sofa. When using open vocabulary, the description could even cover both objects: a sofa covered with a blanket.
Additionally, we qualitatively demonstrate the capability of the benchmark to process open-vocabulary queries. We use the VLMaps’ binary masking method without prompt engineering or dilation as the comparison method. From the results presented in Figure 1, where the matches are shown in orange and non-matches in blue, it can be seen that OpenScene query results are considerably less noisy and encompass entire objects, indicating that the features are more distinctive. While LSeg yields the objects more accurately, the more abstract queries, such as ”striped” or ”relax”, produce no results. OpenScene seems to be more sensitive, finding solutions to abstract queries but yielding many false positive results. While this is in line with the findings of the other experimental results, it might hint towards OpenSeg having a broader vocabulary.
V-C2 Consistency
The intra-map consistency
@@ V-C2 Consistency
The results are presented in Figure 2, in which the intra-map consistency ratios of the methods are illustrated with histograms. Additionally, the means of data are shown with dashed lines. Notably, the ordering of the methods according to their intra-map consistency is the same as in the queryability experiment presented in Section V-C1. OpenScene has better consistency between the labels with each encoder, which is indicated by the lower means of the distributions. This indicates that OpenScene can cluster the embeddings of labels better using the ensemble features. VLMaps with OpenSeg has a long tail that continues beyond the figure, with several outliers.
@@
However, there are multiple classes with a ratio above one. This means that the embeddings in a class are more different from each other than the map on average. This means these classes are challenging to cluster and hard to distinguish. However, these classes have, on average, 2-4% the number of voxels compared to the average number of voxels in the whole data set. For example, the two most represented categories in the tail are ”misc” and ”objects”, both relatively non-descriptive generic categories, and therefore, this could be attributed to the limited number of classes in the experiment.
The inter-map consistency
@@
In Figure 3, the Wasserstein distances between matching, i.e., the label is the same, and non-matching sets of embeddings across all maps are shown. The medians of the data are shown with dashed lines. The ratio between the medians of matching and non-matching labels suggests that OpenScene separates the labels better than VLMaps, and LSeg separates them better than OpenSeg. The ratio of medians of OpenScene with LSeg is 9.21, OpenScene with OpenSeg is 3.67, VLMaps with LSeg 1.65, and VLMaps with OpenSeg 1.24, which is once again the exact ordering observed in Sections V-C1 and V-C2.
@@
While the shapes of distributions are similar, OpenScene has much larger distances on average, even with matching labels. This results from the fact that the distilled features of OpenSeg are of a larger magnitude than the image features of both VLMaps and OpenScene image features. For example, the average norm of all OpenScene with OpenSeg embeddings in the map of the first sequence is 13.68 times larger than the average norm of all VLMaps with OpenSeg embeddings in the same map, and the maximum being 57.97 times larger than the maximum VLMaps feature.
@@
Furthermore, in Figure 4, the distances of matching and non-matching distances are shown per class with orange and blue bars, respectively. The same phenomena can be seen: LSeg distinguishes the classes better than OpenSeg and OpenScene better than VLMaps, which is indicated by the separation of the blue and orange box plots. Additionally, the ten most common and uncommon categories in the Matterport3D data set are marked with red and blue names, respectively. VLMaps performs well in separating common classes and worse with uncommon classes, especially using LSeg, whereas OpenScene performs better across the whole set of labels.
@@
When considering all these results, the combination of OpenScene and LSeg produces maps that are both more queryable and consistent, making it the method that currently can be expected to generalize better and lead to more consistent performance in downstream robotic applications.
VI Conclusion
Mobile robot maps must be able to capture the semantics of the environment for the robots to perform complex tasks. If the semantic representation is constrained to a closed set of classes, the robot’s ability to understand the environment is constrained to the same set. Latent semantic maps present a way to overcome this limitation and enable open-vocabulary scene understanding by representing the semantics in a latent space with a notion of distance.
@@ VI Conclusion
In this work, we propose quantitative metrics to evaluate the quality of state-of-the-art latent semantic maps. In our experiments, OpenScene mapping using 3D scene structure with LSeg semantic embedding performed the best. This indicates that the 3D structure of the environment is useful for inferring the latent semantics. While LSeg performed best in the semantic segmentation setting, the qualitative demonstration hints that OpenSeg might be more sensitive to rarer and more abstract queries when properly thresholded.
The ability of visual-language maps to capture latent semantics is, however, still limited. Currently, fully supervised semantic segmentation methods outperform all latent semantic maps. The relatively low F1-scores and IoUs, especially in the instance-based test, indicate that there is still work to be done to address the problems of thresholding and clustering the query results to segment object instances.
A data set providing ground truth open-vocabulary semantics is direly needed. This would allow extending the evaluation presented here to adequately address the breadth of vocabulary and capture the full promise of VLMs.

### v2

@@ V Experiments
The three main questions that we aim to answer with the experiments are:
@@
    How distinct are their visual-language embeddings within and across maps?
  3. 3.
    How does the grid map resolution affect the queryability?
@@ V-A Data set
The data set used in the study is the Matterport3D data set [31], which consists of houses captured with an RGBD camera and ground truth semantic segmentation. The data set was used in the original work of VLMaps, in which they selected a sample of 10 sequences from the data set. In this work, we use the same sequences and the same set of 40 labels [31] forming the set $\mathcal{L}$; their names are also used in the following experiments as the query set $\mathcal{Q}$ to leverage the semantic ground truth provided with the dataset [^2] Supplementary material including the list of labels will be available in arXiv and link provided here upon acceptance of this work..
@@ V-B Parametrization of the methods
Instead of using the 2D mapping method proposed in the original work [4], the VLMaps map was created using the new 3D mapping method available at the repository of the original authors [33]. This choice was made to create a fair comparison using state-of-the-art methods. The central idea is the same, except the 3D map avoids the problems when aggregating the embeddings along the $z$-axis. All of the parameters were the default used by the original authors.
@@
OpenScene maps were created using the parameters proposed in the original paper. The main analysis uses the proposed 2D-3D ensemble features, but both 2D image features (“fusion”) and 3D point cloud features (“distill”) were created for the ablation study.
@@
V-C Querying the maps
VLMaps proposes a method for performing open-vocabulary queries, which we will refer to as VLMaps querying. When the map is queried with a query $q$, the embeddings on the map are compared using cosine similarity to two embeddings, embedding of $q$ and embedding of string “other”. The binary classification mask is $true$ for voxels where the most similar query is $q$; otherwise, it is $false$. The binary mask is then post-processed by a sequence of binary closing, gaussian blurring, thresholding, and binary dilation to encompass whole object instances.
OpenScene does not propose a method to perform open-vocabulary queries. In the original work, the map is evaluated in a semantic segmentation setting, and open-vocabulary queries are evaluated qualitatively. In our experiment, we implement the VLMaps querying method on OpenScene and use the same parametrization for both methods. Additionally, we evaluate both methods in a semantic segmentation setting, where a binary query mask is acquired by comparing the equality of the voxel labels on the map and the query label.
Finally, VLMaps proposed to use prompt engineering by creating a list of queries by setting $q$ and string “other” into a list of phrases such as “A photo of …” and “there is … in the scene”. Then, the augmented lists of prompts are embedded, and the embeddings are averaged to acquire two embeddings, one for $q$, one for “other”, for comparison. We evaluate the effect of prompt engineering on queryability on VLMaps.
V-D Resolution analysis
For real-world applications, where the robot hardware capabilities might be limited, the memory footprint of the map can be crucial. Small cell sizes can lead to prohibitively large maps, and the size of a grid map is a function of the resolution of the map, as shown in Figure 2. OpenScene has smaller map sizes due to downsampling the images. To investigate the effects of cell size on the map quality, all of the experiments were performed with a set of resolutions $r\in[0.02,0.05,0.1,0.2]$ m. The resolutions proposed in the original works are $0.02$m and $0.05$m in OpenScene and VLMaps, respectively. If not explicitly stated otherwise, we show the results in a resolution of $0.02$m for fair comparison.
VI Results
VI-A Queryability
The results of the queryability are presented in Table I, where the F1-score, precision, recall, and IoU of all of the method and encoder combinations are presented for the semantic segmentation and VLMaps querying.
VI-A1 Semantic segmentation
In semantic segmentation, overall, OpenScene performs better than VLMaps with both encoders, based on the higher F1-scores and IoUs. Furthermore, LSeg performs better with both methods than OpenSeg. Therefore, these results show that both the mapping method and choice of encoder matter in the map creation process. They also indicate that the 3D structure of the environment, included by OpenScene, contains information that can be leveraged in addition to the purely image-based creation of embeddings.
Figure 3: The F1-score of the methods in semantic segmentation task as a function of the map cell size.
VI-A2 VLMaps querying
In the queryability experiment using VLMaps querying, both VLMaps methods performed substantially better than OpenScene. While the results between semantic segmentation and VLMaps querying are not directly comparable, overall low F1-scores indicate that comparison with a single label, “other”, is not sufficient. It is unlikely that any one label would suffice as a universal complement for arbitrary open-vocabulary queries to threshold the results accurately. In a semantic segmentation setting, when the label set is suitable to the data set, the larger set of comparison labels could alleviate this problem by allowing more points of comparison to threshold the query results.
VI-A3 Prompt engineering
One approach to improve the matching of text queries proposed in OpenScene and VLMaps is to use prompt engineering. In our experiment, prompt engineering improves the precision of VLMaps methods, yielding an average $0.059$ and $0.008$ increase in the F1-score on LSeg and OpenSeg, respectively. While OpenScene proposes to use a single prompt augmentation “a XX in a scene”, the set of prompt engineering phrases used in VLMaps did not work with OpenScene. This indicates that the encoders are sensitive to prompt engineering and that a single set of phrases might not work with all encoders. Therefore, while prompt engineering could improve the performance, it does not solve the problem of open-vocabulary query thresholding.
VI-A4 Resolution
The effect of map voxel size, i.e., the resolution, to the F1-score is shown in Figure 3. We present only the F1-score for presentation clarity, as the precision/recall ratio remains similar over the resolutions. The effect of resolution change is minimal in VLMaps as image features are naturally more scale-invariant, as objects of the same category are perceived from different distances in the training data. The 3D features used by OpenScene are not scale invariant, as the point cloud measurements are on an absolute scale. This leads the performance to decrease rapidly. As the difference between the training resolution and the map resolution grows, the number of out-of-distribution samples grows. This effect is especially pronounced in OpenScene using LSeg.
VI-A5 Open-vocabulary
Additionally, we qualitatively demonstrate our evaluation metrics with open-vocabulary queries. The queries, which are only used in this qualitative demonstration, are: “sofa”, “couch”, “white sofa”, “striped”, and “relax”. We use the VLMaps’ binary masking method without prompt engineering or dilation as the comparison method. From the results presented in Figure 1, where the matches are shown in orange and non-matches in blue, it can be seen that OpenScene query results are considerably less noisy and encompass entire objects, indicating that the features are more distinctive. While LSeg yields the objects more accurately, the more abstract queries, such as “striped” or “relax”, produce no results. OpenSeg seems to be more sensitive, finding solutions to abstract queries but yielding many false positive results. While this is in line with the findings of the other experimental results, it might hint towards OpenSeg having a broader vocabulary.
@@ VI-A5 Open-vocabulary
Figure 4: The intra-map distinctness ratios of the methods as histograms, with a kernel density estimation, which is smoothed with a Gaussian kernel. The mean of the distribution is depicted with a vertical line of corresponding color.
@@
VI-B Distinctness
VI-B1 Intra-map distinctness
@@
The results are presented in Figure 4, in which the intra-map distinctness ratios of the methods are illustrated with histograms. Additionally, the means of data are shown with dashed lines. LSeg has better distinctness between the labels with each method, which is indicated by the lower means of the distributions. This indicates that LSeg features of semantic classes are more distinct from each other than in OpenSeg. Furthermore, the 3D ensemble features of OpenScene help to make the features slightly more distinct, as OpenScene has lower means than VLMaps with comparable encoders.
More distinct features, with lower intra-map distinctness ratios, should be preferable, as that indicates better queryability performance. The intuition is that in a best-case scenario, where the query is very accurate, i.e., close to the desired map features in the embedding space, tightly clustered classes are retrieved as a whole, which should lead to better results. More spread-out embeddings are more likely to return partial and sporadic results.
@@
There are multiple classes with a ratio above one, which means that the embeddings in those classes are more different from each other than the map on average. This implies that these classes are challenging to cluster and hard to distinguish. While OpenScene with LSeg does not have any class with a ratio over $1$, OpenScene with OpenSeg has five: “fireplace”, “lighting”, “beam”, “lighting”, “shower”. These classes amount to, on average, $0.8\%$ number of voxels in the whole data set and are, therefore, rare tail classes. VLMaps with both encoders have a larger number of classes with a ratio over one. Notably, both include large classes “floor” and “table”, consisting of $4.44\%$ and $1.67\%$ of the data set, respectively. This implies that spatial features could be helpful in distinguishing large background classes, such as “floor”.
VI-B2 Inter-map distinctness
In Figure 5, the Wasserstein distances between matching, i.e., the label is the same, and non-matching sets of embeddings across all maps are shown, i.e., the blue distances show the similarity between embeddings of the same label between different maps, whereas the orange distances are similarity between embeddings of different classes. The medians of the data are shown with dashed lines. Again, a larger separation between matching and non-matching is preferable, as this would imply that classes remain distinct between environments, i.e., a door would look more like other doors than other classes.
@@ VI-B2 Inter-map distinctness
The ratio between the medians of matching and non-matching labels suggests that OpenScene separates the labels better than VLMaps, and LSeg separates them better than OpenSeg. The ratio of medians of OpenScene with LSeg is 6.31, OpenScene with OpenSeg is 3.78, VLMaps with LSeg 2.07, and VLMaps with OpenSeg 1.42, which is the same ordering observed in queryability results in Section VI-A.
@@
OpenScene has larger Wasserstein distances on average, even with matching labels, which results from the fact that the 3D features of OpenSeg are of a larger magnitude than the image features of both VLMaps and OpenScene image features. For example, the average norm of all OpenScene with OpenSeg embeddings in the map of the first sequence is 13.68 times larger than the average norm of all VLMaps with OpenSeg embeddings in the same map, and the maximum being 57.97 times larger than the maximum VLMaps feature.
@@
Furthermore, in Figure 6, matching and non-matching distances are shown per class with orange and blue bars, respectively. The same phenomena can be seen: LSeg distinguishes the classes better than OpenSeg and OpenScene better than VLMaps, which is indicated by the separation of the blue and orange box plots. Additionally, the ten most common and uncommon categories in the Matterport3D data set that are present in each set of predictions are marked with blue and red names, respectively. VLMaps performs well in separating common classes and worse with uncommon classes, especially using LSeg, whereas OpenScene performs better across the whole set of labels.
@@
When considering all these results, the combination of OpenScene and LSeg produces maps that are more queryable and distinct, making it the method that can currently be expected to lead to more accurate performance in downstream robotic applications. However, OpenScene must be used with the resolution that the 3D encoder is trained on, whereas VLMaps can generalize better to different robot hardware, as it can produce maps of different resolutions, allowing the creation of maps with small memory footprint, with no significant reduction in quality.
VII Conclusion
In this work, we evaluate the quality of state-of-the-art latent semantic maps. We argue that success in downstream tasks is not alone sufficient evidence of the quality of the maps, but rather the representations themselves should be evaluated to understand their limitations. Therefore, we propose quantitative metrics to enable a fair comparison between methods.
@@ VII Conclusion
In our experiments, OpenScene mapping using 3D scene structure with LSeg semantic embedding performed the best, which indicates that the 3D structure of the environment should be used for inferring the latent semantics. While LSeg performed best in the semantic segmentation setting, the qualitative demonstration hints that OpenSeg might be more sensitive to rarer and more abstract queries when properly thresholded. Furthermore, while 2D image features are naturally scale-invariant, the 3D features are susceptible to resolution changes. In use cases with limited hardware resources, the image-based methods work with various resolutions, whereas 3D features must be trained for each resolution. This improves the adaptability of image-based methods to various use cases in real-world settings.
The ability of visual-language maps to capture latent semantics is, however, still limited. Currently, fully supervised semantic segmentation methods outperform all latent semantic maps. The open-vocabulary querying method, proposed in VLMaps, comparing the query to the string “other” cannot sufficiently partition the environment, yielding poor results. This means that a practitioner must choose between the accuracy of the fine-tuned methods with more limited semantics and the broad vocabulary of the less accurate open-vocabulary methods. The relatively low F1-scores and IoUs indicate that in open-vocabulary settings, the thresholding and clustering of the query results remain an open problem.
The growing adoption of latent semantic maps in robotics is already allowing more natural human-robot interaction and complex mobile task execution. The results presented in this work provide evidence that rigorously studying the qualities of these representations is important both to inform practitioners integrating these maps in their robotic pipelines and outlining open problems for researchers to explore.